HETAV PATEL

Grade 11,

What is theNo of neutrons in 5 moles of H2O molecule?

What is theNo of neutrons in 5 moles of H2O molecule?

Grade:10

1 Answers

sam
6 years ago
in 1 H2O molecule there are 0 neutrons(from11H) + 8 ( all 8 from 1 oxygen) neutrons = 8 neutrons
thus in 1 mole
No. of NeutronS= N \times 8
Thus in 5 moles, There are 5\times  N \times 8 neutrons.